When yellow light with a wavelength of 589 nm enters from the air into a transparent plastic body at an inlet angle of 20.0 °, the angle of refraction is 13.3 °. Count

  1. refractive index of plastic 
  2. wavelength of light in plastic.
1
Expert's answer
2021-08-16T16:58:38-0400

1)

"n=\frac{\sin\alpha}{\sin\beta}=\frac{\sin 20.0°}{\sin13.3°}=1.487,"

2)

"\lambda=\frac{\lambda_0}n=\frac{589}{1.487}=396~nm."